what is the width of a 116 square feet room with the length of 12 1/4

Answer by unlockmath(1688)

Hello,
We know the area is Length times width. Let w represent width.
12 1/4 w = 116
Change 12 1/4 to:
49/4 w = 116
Now multiply by 4/49 to get:
w = 464/49 or written as:
9 23/49 feet for the width.
Make sense?
RJ
www.math-unlock.com